### v4

# Beyond Instruction Following: Evaluating Inferential Rule Following of Large Language Models

Abstract
Although Large Language Models (LLMs) have demonstrated strong instruction-following ability, they are further supposed to be controlled and guided by inferential rules in real-world scenarios to be safe, accurate, and intelligent. This demands the possession of inferential rule-following capability of LLMs. However, no prior work has made a clear evaluation of the inferential rule-following capability of LLMs. Previous studies that try to evaluate the inferential rule-following capability of LLMs fail to distinguish the inferential rule-following scenarios from the instruction-following scenarios. Therefore, this paper first clarifies the concept of inferential rule-following and proposes a comprehensive benchmark, RuleBench, to evaluate a diversified range of inferential rule-following abilities. Our experimental results on a variety of LLMs show that they are still limited in following rules. Our analysis based on the evaluation results provides insights into the improvements for LLMs toward a better inferential rule-following intelligent agent. We further propose Inferential Rule-Following Tuning (IRFT). The experimental results show that through IRFT, LLMs can learn abstract inferential rule-following abilities from purely synthetic data and then generalize to RuleBench. The data and code can be found at: https://anonymous.4open.science/r/llm-rule-following-B3E3/
Beyond Instruction Following:
Evaluating Inferential Rule Following of Large Language Models
Wangtao Sun1,2, Chenxiang Zhang1, Xueyou Zhang1, Xuanqing Yu3,7, Ziyang Huang1,2, Haotian Xu6, Pei Chen4, Shizhu He1,2, Jun Zhao1,2, Kang Liu1,2,5Corresponding author: Kang Liu (kliu@nlpr.ia.ac.cn) 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3CAS Engineering Laboratory for Intelligent Industrial Vision, Institute of Automation, Chinese Academy of Sciences, Beijing, China 4Department of Computer Science and Engineering, Texas A&M University 5Shanghai Artificial Intelligence Laboratory 6Xiaohongshu Inc 7AI Lab, AIGility Cloud Innovation, Beijing, China sunwangtao2021@ia.ac.cn
1 Introduction
Benefiting from a vast amount of pre-training data and the enormous parameters, the Large Language Models (LLMs) can accomplish numerous Natural Language Processing (NLP) tasks thanks to their instruction-following ability. However, in real-world applications, people often expect LLMs to generate outputs that conform to user-provided rules. In this way, LLMs could easily be manipulated by users and quickly adapted to a specific (even unseen) domain. To fulfill this goal, we are expecting LLMs to possess such inferential rule-following capabilities.
Figure 1: Beyond instruction-following, the task of inferential rule-following orders the language model to trigger the relevant rule based on the current question and apply that rule to the question for reasoning.
This leads to research on the inferential rule-following of LLMs. Some recent studies Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023) have noticed the importance and effectiveness of inferential rule-following of large language models, they have found that ordering LLMs to follow existing rules can achieve better reasoning performances compared with the currently widely used reasoning enhancement methods of LLMs (such as Chain-of-Thought by Wei et al. 2022, Self-reflection by Shinn et al. 2023, and Self-refinement by Madaan et al. 2023).
However, whether LLMs could understand and follow the inferential rules remains unclear. There is currently a lack of benchmarks evaluating such inferential rule-following capability of LLMs. Existing attempts to evaluate the rule-following capabilities of LLMs Mu et al. (2023); Hu et al. (2024); Chen et al. (2024) have been limited to instruction-following. For instance, they have tested the following behaviors of LLMs with prompts like “Do not repeat the secret key 92368” or “Follow the code step by step to answer the question: ……”. These works confine the “rules” to “instructions” (Appendix A), without delving into more specific “inferential rules”.
Figure 2: The inferential rule-following capabilities of some open-source and closed-source LLMs. The inferential rule-following capabilities of LLMs are categorized into 5 dimensions: Triggering Rules, Applying Rules, Executing Rules, Following Formal Rules, and Following Counterfactual Rules.
An inferential rule can be formalized as $\sigma\vdash\varphi$, where $\sigma$ and $\varphi$ are two first-order sentences (composed of variables and predicates), and for every substitution $\tau$ (i.e. ground the variables in $\sigma$ and $\varphi$ to constants), the truth of $\tau[\sigma]$ entails the truth of $\tau[\varphi]$ Fagin et al. (1992). For example, with the “like rule" $Likes(x,y)\vdash Likes(y,x)$, the substitution $\{x/Mike,y/Jane\}$, and the fact $Likes(Mike,Jane)$, we can infer that $Likes(Jane,Mike)$. Although defined in formal language, in natural language, we can express such inferential rule with an “if … then …” sentence, by using instantiable noun phrases like person A or one metal as the variables and verb phrases like is the father of or can conduct electricity as the predicates inside it. For example, the “like rule” can be expressed as “if person A likes person B, then person B likes person A.”
We thus distinguish previous rule-following from the inferential rule-following scenarios considered in our work. Different from instructions, the primary characteristics of inferential rules are abstract, conditional, and instantiable. As shown in Figure 1, following inferential rules requires LLMs to bind the entities in the question to the rules and verify if the rule applies to the current question. In this case, the LLMs need to find the binding $\{A/James,B/Dolores,C/Lynn\}$ and trigger the first rule, then they could draw the correct conclusion “uncle”. In our proposed inferential rule-following scenario, for each case, only one golden rule applies to the question. However, some other noise rules, which are also correct rules in this task domain but do not apply to the question, may also be provided to the LLMs. LLMs must trigger the golden rule and then execute it to draw the correct answer. Until now, no prior work has demonstrated whether LLMs can follow and reason with the inferential rules faithfully.
Figure 3: The different settings evaluated in RuleBench, including rule quantities, rule forms, Chain-of-Thought (CoT) in applying rules, counterfactual rules, and behavior analysis.
Therefore, beyond the instruction-following studies by previous works, this paper evaluates the LLMs’ capability of inferential rule-following in various reasoning tasks within the scope of inferential rules. This paper proposes a rule-following benchmark, RuleBench, for evaluating the inferential rule-following capability of LLMs under multiple inferential rule-following scenarios, including relation extraction, content moderation, commonsense QA, science QA, and judgment prediction.
Based on RuleBench, this paper has evaluated multiple open-source and closed-source LLMs (§4.1) to find out whether LLMs could understand and follow the inferential rules. In specific, we answer the following questions:
  • Are inferential rules helpful to the reasoning of LLMs? (§4.2)
  • Should inferential rules be presented in natural language or formal language? (§4.3)
  • Is LLMs with Chain-of-Thought (CoT) able to effectively apply the rules? (§4.4)
  • Can LLMs still follow inferential rules in counterfactual scenarios? (§4.5)
This paper also analyzes the cases where LLMs fail to follow the rules (§4.6), categorizing them into Triggering Error and Execution Error, which stand for the cases where LLMs fail to trigger the golden rule and LLMs fail to execute the golden rule, respectively.
Based on these results, as shown in Figure 2, we categorize the inferential rule-following capabilities of LLMs into 5 dimensions (§4.7) to help us intuitively grasp the inferential rule-following capabilities of these LLMs.
Finally, to further improve the inferential rule-following capabilities of LLMs, we propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases (§5). The experimental results show that through IRFT, LLMs can learn abstract inferential rule-following abilities from purely synthetic data and then generalize to RuleBench. In summary, the major contributions of this paper are as follows:
  • We introduce inferential rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following.
  • We leverage and re-process the existing reasoning benchmarks and propose an inferential rule-following benchmark, RuleBench, for evaluating the inferential rule-following capability of LLMs.
  • We evaluated the capabilities of inferential rule-following of multiple open-source and closed-source LLMs on various tasks and rule settings, and categorized their inferential rule-following abilities into 5 dimensions. Based on the results, we analyze the possible reasons that limit the inferential rule-following capabilities of current LLMs and provide some insights into the improvements for LLMs toward a better inferential rule-following intelligent agent.
  • We propose the Inferential Rule-Following Tuning (IRFT) that enables LLMs to learn to trigger and apply the correct inferential rule based on the current cases. The experimental results show that through IRFT, LLMs can learn abstract inferential rule-following abilities from purely synthetic data and then generalize to RuleBench.
2 Related Work
2.1 Rule-enhanced LLM Reasoning
While LLMs have demonstrated remarkable zero-shot reasoning capabilities in many downstream tasks, they still generate outputs that do not conform to logic or human preference. Some research studies have found that compared with the reasoning enhancement methods based on LLMs themselves like Chain-of-Thought Wei et al. (2022), Self-reflection Shinn et al. (2023), and Self-refinement Madaan et al. (2023), providing LLMs with relevant rules with Retrieval-Augmented Generation (RAG) paradigm do better in helping them conduct reasoning in the downstream tasks Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023); Zhao et al. (2023). However, the inferential rule-following capability of LLMs is far from satisfactory. No prior work has comprehensively evaluated whether LLMs can benefit from the provided rules under different scenarios and how LLMs can follow the rules better. To make up for this gap, this paper conducted a series of experiments to evaluate the inferential rule-following capabilities of several open-source and closed-source LLMs and provide some insights into how LLMs can follow rules better.
2.2 LLMs Instruction-following
Instruction-following has been generally considered an important capability of LLMs Zhong et al. (2021); Mishra et al. (2021); Wei et al. (2021); Yin et al. (2023), and some previous works have been done to evaluate the instruction-following capability of LLMs Zhou et al. (2023); Qin et al. (2024). However, only a few works have cast their attention to the question of inferential rule-following. Recent works focused on the rule-following capability of LLMs Mu et al. (2023); Hu et al. (2024); Chen et al. (2024) confined the rule-following to instruction-following. Instead, this paper proposes the scenario of inferential rule-following and sets up useful baselines for future works.
3 RuleBench
To construct RuleBench, we have leveraged and re-processed the existing reasoning benchmarks for different inferential rule-following scenarios, including relation extraction (CLUTRR, Sinha et al. 2019), content moderation (SALAD, Li et al. 2024), commonsense QA (DEER, Yang et al. 2022 and ULogic, Wang et al. 2024), mathematics QA (TheoremQa, Chen et al. 2023), and judgment prediction (CAIL2018, Xiao et al. 2018; Zhong et al. 2018). The details of the construction of each benchmark and the prompts used during constructing RuleBench can be found in Appendix B.
Under the scenarios introduced above, As shown in Figure 3, RuleBench involves multiple settings of inferential rule-following, to comprehensively evaluate the LLMs from different perspectives. The settings include rule quantity (i.e. how many rules are provided to the LLMs while only one of them applies to the current case), rule form (i.e. which form the rules illustrated in, natural language or formal language), the presence of Chain-of-Thought when applying rules (i.e. directly generate the answer based on the question and rules, or trying verbally apply the rule to the question before answering it), and rule factuality (i.e. whether the conclusion of the rule is factual or counterfactual). RuleBench allows us to analyze the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error (i.e. LLMs fail to trigger the golden rule) and Execution Error (i.e. LLMs success to trigger the golden rule but fail to execute the golden rule).
4 Evaluation
To comprehensively evaluate the inferential rule-following capabilities of LLMs, based on the proposed RuleBench, this paper has designed 5 main parts of experiments. We evaluate the effects of rule quantity (§4.2), rule form (§4.3), the presence of CoT when applying rules (§4.4), and rule factuality (§4.5). Besides, we analyzed the failure cases of inferential rule-following from a behavioral perspective, classifying them into Triggering Error and Execution Error4.6). Based on these evaluation results, we categorize the inferential rule-following capabilities into 5 dimensions and compare the performances of 7 open-source and closed-source LLMs (§4.7). The test-time prompts used in this section can be found in Appendix C.
4.1 Model Selections
For open-source LLMs, we adopt Llama-2-7b-chat Touvron et al. (2023), Meta-Llama-3-8B AI@Meta (2024), Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and Yi Young et al. (2024). For closed-source LLMs, we adopt gpt-3.5-turbo, gpt-4-turbo OpenAI (2023), and gpt-4o from OpenAI. The comprehensive performance comparison of them is shown in Figure 2 and the explanation and analysis is in §4.7.
Figure 4: The inferential rule-following performance of LLMs under different rule quantities.
Model Form CAIL2018 SALAD TheoremQA CLUTRR ULogic DEER
All Few All Few All Few All Few All Few All Few
gpt-4o FOL 95.78 95.18 92.93 96.20 70.36 69.45 43.13 43.10 87.59 70.60 73.81 38.10
NL 96.99 97.59 95.19 95.35 69.45 69.45 58.97 62.67 94.82 79.16 92.86 71.43
Llama-2-7b-chat-hf FOL 50.60 53.01 58.27 81.22 39.16 38.18 23.00 27.48 60.36 79.04 57.14 71.43
NL 53.61 63.25 72.59 88.91 39.00 35.96 24.33 40.36 60.00 83.86 59.52 59.52
gpt-3.5-turbo FOL 84.94 86.14 73.33 96.83 49.82 53.26 41.54 53.28 63.73 66.10 71.43 78.57
NL 90.36 89.76 80.64 99.24 50.18 50.36 42.32 59.36 80.72 70.48 78.57 83.33
Yi-1.5-6B-Chat FOL 96.99 97.59 64.36 70.28 53.27 52.00 23.95 23.47 90.12 81.93 64.29 52.38
NL 96.99 95.78 75.73 80.44 52.00 49.64 17.27 21.37 87.59 82.77 73.81 66.67
gpt-4-turbo FOL 96.39 95.78 93.25 93.72 66.55 65.09 46.95 53.99 98.43 98.55 90.48 88.10
NL 96.99 98.80 94.81 94.70 66.36 64.36 49.32 53.92 98.43 98.67 88.10 90.48
Meta-Llama-3-8B FOL 89.16 93.98 77.71 77.16 48.91 50.73 38.55 41.89 92.17 97.23 83.33 80.95
NL 89.76 97.59 82.28 77.92 48.18 48.55 37.40 45.52 92.17 95.06 85.71 85.71
Mistral-7B-Instruct FOL 78.31 79.52 72.62 80.54 17.45 21.45 25.10 29.20 86.75 92.41 64.29 64.29
NL 84.34 92.77 82.36 83.43 22.18 24.55 26.62 32.73 95.42 94.46 57.14 64.29
Table 1: The of LLMs on RuleBench with rules of formal language (FOL) and natural language (NL).
Figure 5: The inferential rule-following performance of LLMs when applying rules with or without using Chain-of-Thought. The dashed box indicates the improvement (positive or negative) of w/ CoT over w/o CoT.
4.2 Inferential Rules Are Helpful for the Reasoning of LLMs
To evaluate whether inferential rules are helpful for the reasoning of LLMs, we adopt the following settings to test the LLMs.
  • No Rule. This setting simply prompts the LLMs with the original question and without the inferential rules.
  • Golden Rule. This setting prompts the LLMs with the golden rule (i.e. an inferential rule that should be applied to the question) together with the original question.
  • Few Rule. This setting prompts the LLMs with the golden rule and two random noise rules together with the original question.
  • All Rule. This setting is similar to Few Rule while the number of noise rules increases to 30. This setting simulates a scenario where users prompt the LLMs with all possible inferential rules in the tasks instead of the relevant rules retrieved based on the query.
All these rule settings are tested in a zero-shot manner. As shown in Figure 4, in most cases, LLMs enjoy great performance improvements while being prompted with one golden inferential rule (No Rule $\rightarrow$ Golden Rule). Nevertheless, as the number of noise rules increases, LLMs will find it hard to trigger and leverage the golden rule and thus have a performance drop (Golden Rule $\rightarrow$ Few Rule $\rightarrow$ All Rule).
Besides, we find that by following inferential rules, LLMs have better performance improvements on tasks that require complex reasoning, such as CLUTRR and CAIL2018. On the commonsense reasoning tasks, as the LLMs have parametric knowledge, the performance improvements brought by following inferential rules are relatively slim. Moreover, we find that all LLMs fail to follow the inferential rules in TheoremQA, which illustrates the defect of current LLMs that can not follow complex mathematical or physical rules.
4.3 LLMs Prefer Natural Language Rules than Formal Language Rules
Formal language is widely used in early Artificial Intelligence, which is able to conduct efficient and generalized reasoning. However, LLMs have shown competitive or even superior reasoning performance over traditional formal language rule-based engines, i.e. Knowledge Graphs Luo et al. (2023). In contrast to formal language rule-based reasoning, reasoning with LLMs is more flexible and robust to various data and tasks. Therefore, we would like to know if we can combine these two paradigms, i.e. whether LLMs can follow formal language rules.
To evaluate whether LLMs can follow formal language rules, we transform the natural language rules of each benchmark into the form of First-Order Logic (FOL) by executing deterministic functions or prompting ChatGPT (Appendix B). Then we compare the reasoning performances of LLMs which are prompted by different forms of inferential rules in both All Rule and Few Rule settings.
As shown in Table 1, in most cases, LLMs conduct reasoning better with natural language rules than formal language rules (except for TheoremQA and Yi-1.5-6B-chat). This aligns with our intuition that LLMs are mostly pre-trained with natural language and thus the inferential rules expressed with natural language are closer to the pre-trained distributions of LLMs than the inferential rules expressed with formal language. This confirms the research motivation and value of Symbol-LLM Xu et al. (2023). Nevertheless, in most cases, LLMs still can follow the formal language rules. This reveals the possibility of learning formal language rules from a symbolic reasoning engine and then using LLMs for neural inference.
Model Factuality CAIL2018 SALAD CLUTRR ULogic
Golden Few Golden Few Golden Few Golden Few
gpt-4o Factual 98.19 96.99 99.74 97.93 81.18 65.46 88.71 96.55
C.F. 97.14 97.14  8.22 68.42 37.50  5.53 86.87 88.80
Llama-2-7b-chat Factual 62.65 70.63 70.68 41.54 77.58 37.79 91.81 87.71
C.F. 79.25 67.21 26.06 52.12 71.37 13.93 67.35 53.61
Yi-1.5-6B-Chat Factual 98.19 95.18 87.15 66.27 78.72 37.02 98.55 97.71
C.F. 87.43 80.57 71.24 56.68 70.13 19.27 73.61 71.08
gpt-3.5-turbo Factual 99.40 98.19 99.74 99.48 95.49 63.83 97.11 96.75
C.F. 97.14 91.43 100.0 100.0 98.50 43.86 81.69 75.78
gpt-4-turbo Factual 99.40 98.19 100.0 97.98 82.70 63.49 98.58 98.31
C.F. 99.43 100.0 12.93 77.55 78.72 36.16 86.51 86.99
Meta-Llama-3-8B Factual 89.16 86.14 96.09 93.35 62.21 40.17 98.92 98.80
C.F. 62.29 86.29 100.0 94.00 39.50  6.49 79.64 75.90
Mistral-7B-Instruct Factual 98.19 88.55 90.88 88.19 50.38 35.88 97.17 93.36
C.F. 88.00 65.14 62.10 87.48 26.15 13.26 33.61 24.10
Table 2: The performance of LLMs on RuleBench when following factual and counterfactual (C.F.) rules.
4.4 Chain of Thought Is Inadequate for LLMs to Apply Inferential Rules
Chain-of-Thought Wei et al. (2022) has been widely verified as a useful prompting technique to help LLMs conduct multi-hop reasoning. To evaluate whether LLMs can use CoT to apply inferential rules in the inferential rule-following scenario, we choose the few-shot Golden Rule and Few Rule settings. We created two demonstrations with CoT and two demonstrations without CoT under such settings for LLMs to conduct In-context Learning.
However, as shown in Figure 5, LLMs with CoT have not exhibited stronger inferential rule-following performances in most cases. This may be attributed to the lack of planning of CoT. CoT conducts straightforward reasoning from the question to the answer with multiple reasoning hops. However, when applying the inferential rules, it involves trying to apply each rule to the current question and thinking about whether to execute this rule. Therefore, plain CoT is inadequate for LLMs to apply the inferential rules. Prompting techniques (e.g. Tree of Thought, Yao et al. 2024) or decoding algorithms (e.g. KCTS, Choi et al. 2023) that involve planning steps are needed for helping LLMs to apply the inferential rules.
4.5 LLMs Struggle to Follow Counterfactual Inferential Rules
Although we have verified the effectiveness of the inferential rules, it is still unclear whether LLMs strictly follow the given inferential rules or merely the rules activate their parametric knowledge. Therefore, we designed the scenario of counterfactual rule-following.
To evaluate whether LLMs can follow counterfactual rules, we construct corresponding counterfactual benchmarks and rule sets of CLUTRR, SALAD, ULogic, and CAIL2018. Specifically, we replace the ground truth of each question and the conclusion of the corresponding rule with a random incorrect answer. So in this counterfactual setting, the LLMs are supposed to generate the “incorrect answer” based on the given counterfactual rules.
As shown in Table 2, in most cases of both Golden Rule and Few Rule settings, LLMs have significant performance drops when following counterfactual rules, compared with following factual rules. These results indicate that the performance improvements brought by following rules are partly attributed to the parametric knowledge of LLMs, besides following inferential rules.
4.6 Behavioral Analysis of LLMs Following Inferential Rules
To understand why LLMs fail to follow the given inferential rules in the reasoning process, we made a behavioral analysis of LLMs in the failure cases of LLMs inferential rule-following. Specifically, we adopt the few-shot Few Rule settings for LLMs to follow the rule-applying demonstrations to apply the given inferential rules to the current question. We ordered the LLMs first to choose an inferential rule to follow and then reason with it. By parsing the output of LLMs we can classify the failure cases of LLMs inferential rule-following into two categories: Triggering Error and Execution Error. Triggering Error indicates that the LLMs choose a noise rule for the current case and therefore lead to an incorrect reasoning result. Execution Error indicates that although LLMs have chosen the correct rule for the current case, they fail to draw the correct conclusion of rule body. To faithfully describe the inferential rule-following behavior of LLMs instead of being affected by the parametric knowledge of LLMs, we run the analysis under the counterfactual settings of the selected benchmarks.
From the results shown in Figure 6, we can tell that when tackling different tasks, LLMs exhibit different behaviors in following rules. While rules have a heavy head for triggering (e.g. in CLUTRR and CAIL2018, the rule head will be a series of relational hops among characters), the LLMs are likely to make Triggering Errors. While the rule head is easy and commonsensical (e.g. in SALAD and ULogic), but the conclusion of the rule body is ambiguous or confused (the counterfactual scenario), the LLMs are likely to make Execution Errors. A case study of these two types of failures can be found in Appendix E.
To avoid Triggering Errors in the scenario of rule-enhanced reasoning with RAG paradigm (§2.1), the rule retriever plays a crucial role. The Triggering Errors can be eliminated if the rule retriever only retrieved the golden rules. However, existing works mostly employ simple sparse retrievers such as BM25 Yang et al. (2023); Zhu et al. (2023), which greatly compromises the inferential rule-following performance of LLMs.
To avoid Execution Errors in following rules, the LLMs need to faithfully execute the rule body and avoid generating conclusions of illusions. Therefore, users may avoid letting LLMs follow the rules that are counterfactual or out of the pre-trained distribution of LLMs before they fine-tune the LLMs to adapt to those domains or specific tasks.
Figure 6: The proportion of Triggering Error and Execution Error produced by LLMs on RuleBench.
4.7 Inferential Rule-Following Capabilities of LLMs
To make a comprehensive evaluation of the inferential rule-following capability of the LLMs, we categorize the experimental results in the previous sections into 5 dimensions: Executing Rules, Triggering Rules, Following Formal Rules, Applying Rules, and Following Counterfactual Rules. The details of these dimensions are shown in Appendix F.
As shown in Figure 2, while the closed-source LLMs show dominant performances in the scenario of inferential rule-following, some open-source LLMs, like Llama-3-8B, exhibit competitive performances and have balanced capabilities in all dimensions. Among the closed-source LLMs, gpt-4-turbo is more capable of following formal language rules while gpt-3.5-turbo shows a stronger capability of following counterfactual rules.
Generally, LLMs are not very good at inferential rule-following. This may be attributed to the lack of training in inferential rule-following in the current LLMs. In the next section, we propose a fine-tuning method to effectively further improve the inferential rule-following capabilities of LLMs.
Dataset CLUTRR SALAD CAIL2018 base 40.36 88.91 63.25 IRFT 42.94 91.80 67.44 Dataset TheoremQA ULogic DEER base 35.52 79.64 52.38 IRFT 43.03 87.83 61.90
Table 3: The performances of base and fine-tuned Llama-2-7b-chat on all datasets of RuleBench. The fine-tuning data is constructed using StringGame’s OOD synthetic data. Tested in zero-shot Few Rule setting.
5 Inferential Rule-Following Tuning
To further improve the inferential rule-following capabilities of LLMs, we propose Inferential inferential rule-following Tuning (IRFT). IRFT involves a golden rule with a few randomly sampled noise rules in the prompt and therefore orders the LLMs to learn to trigger and execute the golden rule. The tuning objective can be formalized as:
$J_{IRFT}=\mathbb{E}_{\begin{subarray}{c}q,r,a\sim p_{train}\\ r_{1},...,r_{n}\sim U(R)\end{subarray}}-\log p(a|[q;r;r_{1},...,r_{n}])$
Where the $q,r,a$ stands for the question, the golden rule, and the answer from the training set, respectively. $r_{i}\sim U(R)$ stands for randomly sampling $n$ rules from the entire rule sets as the noise rules.
To thoroughly separate the rule-following ability from domain knowledge, we propose to use purely synthetic data to construct the corpus, StringGame, for IRFT. The details of StringGame can be found in Appendix G.
As shown in Table 3, after training with IRFT on StringGame, the LLM enjoys a performance improvement on all tasks of RuleBench. We also evaluate both the base LLM and tuned LLM on an instruction-following benchmark, InfoBench, where their accuracies are 74.36% and 74.49%, respectively. Based on these results, we believe that inferential rule-following is an abstract and fundamental capability. Through IRFT, this capability can be abstracted and learned from purely synthetic symbolic tasks, allowing generalization to real-world rule-following tasks. Meanwhile, IRFT does not affect their general instruction-following ability. We also test using in-domain data for IRFT (Appendix H), and the performance improvements brought by IRFT are more significant.
6 Conclusion
In this paper, we introduce inferential rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following. We then construct and propose a new benchmark, RuleBench, for evaluating the inferential rule-following capabilities of LLMs. Based on RuleBench, we conduct a series of experiments to evaluate the inferential rule-following capabilities of 7 open-source and closed-source LLMs from different perspectives. We categorize the inferential rule-following capability in 5 dimensions and provide some insights into improvements for LLMs toward a better inferential rule-following intelligent agent. Finally, we propose the Inferential Rule-Following Tuning (IRFT), which further improves the inferential rule-following capabilities of LLMs.
Limitations
Although IRFT has shown remarkable performances on RuleBench, beyond using IRFT on specific downstream tasks, we are looking forward to extending IRFT to the pre-training stage of LLMs (like IFT), such that it is possible to enable LLMs to master more basic and generalized inferential rule-following capabilities.
Ethics Statement
Our research aims to evaluate the inferential-inferential rule-following capability of LLMs. To mitigate risks associated with some sensitive content in the benchmark, we restrict access to authorized researchers who adhere to strict ethical guidelines. These measures safeguard research integrity while minimizing potential harm.
Appendix A Instructions vs Rules
Nevertheless, we can not confine rules to instructions, or even identify rules with instructions Ribes-Inesta (2000). Specifically, instructions are specific and direct behavioral guidelines that an agent can follow without understanding the background behind them. Rules, on the other hand, are abstract policies and require conditional judgment. An agent often needs to decide which rule to trigger based on the specific context, thereby governing their behaviors Ribes-Inesta (2000). Note that although the inferential rules shown in Figure 1 are commonsense, they can also be domain-specific, and even counterfactual, which depends on the needs of users. Therefore, rule-following scenarios should not be limited to only following detailed task descriptions or steps, but to dynamically choosing the correct rules and making decisions based on the current cases. Following (Fagin et al., 1992), we call this type of rule inferential rule and named the scenario considered in this paper LLM inferential rule-following.
Appendix B Details of Constructing RuleBench
Here are the details of constructing each benchmark in RuleBench. The prompts used in this process are shown in Figure 10,11,12,13,14.
  • CLUTRR Sinha et al. (2019). Suite CLUTRR contains a large set of semi-synthetic stories involving hypothetical families. Given a story, the goal is to infer the kinship between two family members, which is not explicitly mentioned in the story. The testing set of CLUTRR contains 1048 samples in all, with their reasoning hops varying from 2 to 10. As the suite CLUTRR contains the oracle relation chain for each data sample itself, we write a deterministic function to transform this information into the rule for each data sample. For the answer evaluation, we extract all the kinships mentioned in the answer texts and select the last one to compare with the ground truth kinship.
  • SALAD Li et al. (2024). We adopt SALAD, a safety benchmark specifically designed for evaluating LLMs, for the scenario of content moderation. Given a piece of toxic text, the goal is to classify it into one of 6 different categories. The testing set of SALAD contains 5939 samples in all. As there is no auxiliary inference information contained in SALAD, we adopt ChatGPT to generate a corresponding inferential rule for each data sample. Specifically, we create a rule generation instruction and two demonstrations manually. They are prompted to ChatGPT together with each sample in SALAD. Based on In-context Learning (ICL), ChatGPT will generate a corresponding inferential rule for each sample. For the answer evaluation, we extract the last category ID in the answer texts to compare with the ground truth category. Note that, as SALAD involves identifying toxic content, the safety-aligned LLMs will probably refuse to answer the question (Despite the questions of the SALAD being to have LLMs classify toxic content, rather than inducing them to generate toxic content). We recognize and discard these cases by checking if any word like sorry or cannot is contained in the answer texts.
  • DEER Yang et al. (2022). DEER is proposed as a 1.2k rule-fact pairs dataset, about natural and social sciences. Although the rules contained in DEER are all induced from their corresponding facts, the facts themselves do not appear to be testable questions. Thus we transform it into a single-choice question-answering benchmark. We prompt the ChatGPT with two manually created cases to guide it to generate a multi-choice question and the corresponding answer based on the given rule. All question-answer pairs are then verified by humans. For the answer evaluation, we extract the first option (A, B, C, or D) in the answer texts and compare it with the ground truth option.
  • TheoremQA Chen et al. (2023). TheoremQA is a mathematics problem dataset, characterized by the fact that each question and answer has a corresponding theorem. TheoremQA comprises 800 QA pairs covering 350+ theorems spanning across Math, EE&CS, Physics, and Finance. In this dataset, each math problem is associated with a corresponding theorem, but the theorems are not strict inferential rules. Therefore, we used gpt-4-turbo to transform each theorem into an “if … then …” rule format. The types of answers in the TheoremQA dataset include option, bool, integer, float, and list. Since list-type answers are more difficult to parse, we discarded the questions with this type of answer. Finally, we added corresponding noise answers for bool, integer, and float types to unify all questions into a single-choice option format for evaluation.
  • ULogic Wang et al. (2024). Ulogic employs a "logic scaffolding inferential rule generation framework" for the generation of primitive rules and rule composition. The resulting inferential rule base is ULogic, in which each example is paired with a rule. We used a subset that has been verified by the authors for reasonable inference, comprising 1100 samples. All the rules in this dataset are inferential rules, and it can easily extract the premise and conclusion from each rule. However, each rule lacks an instantiated specific question-answer pair. Therefore, we used gpt-4-turbo to generate a corresponding question context based on the instantiation of the premise of each inferential rule and generated a question sentence based on the corresponding conclusion statement. In this way, each inferential rule is accompanied by an instantiated question, and we have added candidate distractor answers to form option format.
  • CAIL2018 Xiao et al. (2018); Zhong et al. (2018). Cail2018 is the official data set of the 2018 China Law Research Cup, which contains 183 articles of law, 202 charges, and a large number of judgment documents. Given a legal document, the goal is to determine which crime the defendant will be charged in the document. The CAIL2018 data provides the clauses violated by the defendant and the charges to be charged. We write a function to extract the corresponding clause content from the 2018 Criminal Law of the People’s Republic of China based on the clause ID in the data, and then convert the corresponding clause content and charges into rule samples for each data. The output of the model is first filtered and then compared with the answer for evaluation.
Figure 7: An example of Triggering Error on CLUTRR.
Figure 8: An example of Execution Error on CLUTRR.
Appendix C Test-time Prompts
The prompts used during test-time are shown in Figure 15,16,17,18,19,20. Please note that golden rules don’t always appear first; they are mixed with noise rules in random order during testing. The orders in these figures are for illustrative purposes.
Appendix D Evaluation of Inferential Rule-following on Long-tail Instances
To verify if following inferential rules is more beneficial when encountering long-tail instances, we selected instances with labels that are among the least frequent labels while ensuring that their amount is no less than 10% of the total dataset in SALAD (As their labels are meaningful) to serve as long-tail instances.
As shown in Table 4, closed-source LLMs (gpt-4o, gpt-4-turbo, gpt-3.5-turbo), compared with the performance on overall instances, although they have performance drops when following only instructions (No Rule) on long-tail instances, still achieve competitive performance while following inferential rules (All Rule, Few Rule, Golden Rule). This indicates that for closed-source LLMs, the improvement brought by following inferential rules on long-tail instances is greater compared to regular instances. However, for open-source LLMs, we found that they experience significant performance drops when following only instructions, and strangely, the effectiveness decreases with an increase in the number of noise rules, opposite to their behavior on the overall instances (Section 4.2). This indicates that these open-source LLMs have poor inferential rule-following capabilities on the distributions of long-tail samples.
Appendix E Case Study of Behavioral Analysis
Dataset Setting gpt-4o gpt-4-turbo gpt-3.5-turbo Llama-2-7b-chat Meta-Llama-3-8B Mistral-7B-Instruct Yi-1.5-6B-Chat
SALAD No Rule 47.25 54.85 53.70 20.12 39.55 53.51 50.94
All Rule 95.19 94.81 80.64 72.59 82.28 82.36 75.73
Few Rule 95.35 94.70 99.24 88.91 77.92 83.43 80.44
Golden Rule 98.59 99.36 99.67 99.67 95.11 92.28 81.78
SALAD long-tail No Rule 40.90 39.97 23.15  7.10  6.64  7.10  6.94
All Rule 95.83 92.28 77.47 75.46 75.15 75.00 74.07
Few Rule 96.91 97.68 99.23 55.71 56.26 56.94 53.70
Golden Rule 98.92 99.69 99.69 28.64 28.95 29.06 29.41
Table 4: The inferential rule-following performance of LLMs on all instances and only long-tail instances of SALAD.
To better illustrate what happens when LLMs fail to follow the inferential rules, we show two cases in which LLMs experience the Triggering Error (Figure 7) and Execution Error (Figure 8), respectively.
In the case of Triggering Error, LLMs appear to trigger an noise rule and consequently draw the wrong answer. In the case of Execution Error, although LLMs have triggered the golden rule, however, they fail to apply the rule to the question correctly, therefore also drawing the wrong answer.
Appendix F Details of The Dimensions
  • Executing Rules. We average the results in all Golden Rule settings to obtain the capability of Execution Rules of LLMs. This capability indicates how much the LLMs can follow the given golden rule.
  • Triggering Rules. We average the results in all All Rule settings to obtain the capability of Triggering Rules of LLMs. This capability indicates how much the LLMs can resist the interruption of noise rules and find the golden rule.
  • Following Formal Rules. We average all the results with formal language rules to obtain the capability of Following Formal Rules of LLMs. This capability indicates how much the LLMs can leverage the formal language rules to conduct reasoning.
  • Applying Rules. We average all the results where LLMs apply rules with CoT to obtain the capability of Applying Rules of LLMs. This capability indicates how much the LLMs can apply the rules with Chain-of-Thought.
  • Following Counterfactual Rules. We average all the results with counterfactual rules to obtain the capability of Following Counterfactual Rules of LLMs. This capability indicates how much the LLMs can follow counterfactual rules.
Appendix G Details of The StringGame
The StringGame aims to use symbolic execution to construct a purely synthetic rule-following training dataset. It involves a series of five key steps that encompass rule generation, rule implementation, rule execution, prompt construction, and CoT generation. This task is about triggering and following the golden rule among the given rules to make the correct calculations based on the given string. An example is shown in Figure 9.
Example of StringGame Input: YTNRCMTZTVEAVQEHVJHW Rules: 1. If the length of the string is even, then output the number of unique consonants in the string. 2. If the number of "G"s is equal to 3, then output the index of the first "G" (index begins from 0). 3. If the string starts with a vowel, then output the total number of letters in the string divided by 2. 4. If the string starts and ends with the same letter, then output the index (1-based) of the first occurrence of the letter "R". Output: Based on the first rule, the answer is 12.
Figure 9: An example of StringGame.
G.1 Rule Generation
In this phase, a rule base is first initialized with 8 manually created seed rules. While the size of the rule base is smaller than 1000, we iteratively sample 5 rules from the rule base as few-shot demonstrations and ask the LLM to generate more similar but diverse rules. The response of the LLM will then be parsed and the new rules will be appended to the rule base.
G.2 Rule Implementation
Then, with two rule-function demonstrations, we use the LLM to implement each rule in the rule base to a snippet of the Python function. After that, another LLM, serving as the judge, will verify whether the generated code matches the provided rule. If the generated code passes the LLM validation and Python syntax check, it will then be saved as rule-function pairs.
G.3 Rule Execution
After all rules have been implemented as functions, we randomly generate strings with upper-case letters as the input. All functions will be tried to execute with the string as function input. If the function successfully returns a number, the corresponding rule will be noted as a “golden rule” for this string (meanwhile the function output will be saved as the answer to the question), otherwise (return None or get a run-time error), this rule will be noted as a “noise rule” for this string.
G.4 Prompt Construction
Having obtained the input strings together with their golden rules and noise rules, for each string, we sample one golden rule and randomly 0-3 noise rules. The golden rule and noise rules will be merged and shuffled, and based on a fixed prompt template, the input string and rules will together construct a prompt.
G.5 CoT Generation
Finally, we leverage STaR Zelikman et al. (2022) to sample success CoT trajectories from a closed-source LLM, i.e. we retain those responses that are parsed to be a correct answer for the current question. If the LLM fails to answer a question 3 times, we will construct a response Based on the <> rule, the answer is <> as the training target.
Rule Setting Dataset CLUTRR SALAD CAIL2018 Golden Rule base 55.25 99.67 78.92 IRFT 100.0 99.90 87.95 Few Rule base 40.36 88.91 63.25 IRFT 89.50 99.00 90.96
Table 5: The performances of base and fine-tuned Llama-2-7b-chat on three datasets of RuleBench. The fine-tuning data is constructed using RuleBench’s in-domain training data.
Appendix H In-domain IRFT on RuleBench
We also try to leverage the in-domain data for IRFT. Based on the training data in RuleBench (only the CLUTRR, SALAD, and CAIL2018 datasets contain training data), we constructed inferential-rule following samples under Few Rule ($n=2$) and Golden Rule ($n=0$) settings for IRFT.
As shown in Table 5, our proposed IRFT further significantly improves the performances of LLMs in the inferential rule-following scenarios and greatly outperforms the base model in both Few Rule and Golden Rule settings.
Figure 10: The prompt used for constructing SALAD.
Figure 11: The prompt used for constructing DEER.
Figure 12: The prompt used for constructing TheoremQA.
Figure 13: The prompt used for constructing ULogic.
Figure 14: The prompt used for constructing CAIL2018.
Figure 15: The test-time prompt used for CLUTRR.
Figure 16: The test-time prompt used for SALAD.
Figure 17: The test-time prompt used for DEER.
Figure 18: The test-time prompt used for TheoremQA.
Figure 19: The test-time prompt used for ULogic.
Figure 20: The test-time prompt used for CAIL2018.